How do you write the slope-intercept equation for a line shown on a graph when the slope is going downward? The points are (4,-1) (3,-3).
Answer by rfer(16322)

slope m=(-3--1)/(3-4)
m=-2/-1
m=2
--------
This is a positive slope
------------------
y=mx+b
-3=2*3+b
-3=6+b
b=-9
-----------
y=2x-9
---------
I guess I don't understand your question based on the information given?